# MSVCOD:A Large-Scale Multi-Scene Dataset for Video Camouflage Object Detection

### v1

@@ Abstract
Video Camouflaged Object Detection (VCOD) is a challenging task which aims to identify objects that seamlessly concealed within the background in videos. The dynamic properties of video enable detection of camouflaged objects through motion cues or varied perspectives. Previous VCOD datasets primarily contain animal objects, limiting the scope of research to wildlife scenarios. However, the applications of VCOD extend beyond wildlife and have significant implications in security, art, and medical fields. Addressing this problem, we construct a new large-scale multi-domain VCOD dataset MSVCOD. To achieve high-quality annotations, we design a semi-automatic iterative annotation pipeline that reduces costs while maintaining annotation accuracy. Our MSVCOD is the largest VCOD dataset to date, introducing multiple object categories including human, animal, medical, and vehicle objects for the first time, while also expanding background diversity across various environments. This expanded scope increases the practical applicability of the VCOD task in camouflaged object detection. Alongside this dataset, we introduce a one-steam video camouflage object detection model that performs both feature extraction and information fusion without additional motion feature fusion modules. Our framework achieves state-of-the-art results on the existing VCOD animal dataset and the proposed MSVCOD. The dataset and code will be made publicly available.
@@ 1 Introduction
Camouflaged Object Detection (COD) [9] focuses on identifying and segmenting the hidden objects that closely resemble their backgrounds. Those camouflaged objects exhibit intricate visual patterns, edge breaking, texture similarities, and color matching, making them blend seamlessly into their surroundings and challenging to detect compared to traditional object detection methods [40, 54]. COD has applications in diverse fields such as medical image segmentation [11, 10], enemy detection in the battlefield [28], art [14, 3], industry [50], scientific research [37], and agriculture [38]. Given that motion and multi-angle visual information can help to detect camouflaged objects, Video Camouflage Object Detection (VCOD) was developed. VCOD, a sub-domain of COD, focuses on segmenting out camouflage targets in the videos by utilizing motion and multi-angle visual cues.
@@
Current video camouflage datasets primarily focus on animal scenes, overlooking the diverse camouflage modalities human society. This limitation restricts the broader applicability of VCOD, particularly in fields such as medicine, security, search-and-rescue, and art. The lack of diverse datasets results in researchers not having enough data to train and test models on benchmark datasets. To enhance model generalization across scenes and objects, existing VCOD models [25, 53] are often pretrained on static image datasets [9]. To address this gap, we built a novel, large-scale multi-scene VCOD benchmark dataset. It includes 162 video clips across four object categories (human, animal, medical, and vehicle) and seven scenarios (aquatic, field, medical, art, jungle, desert, snowfield). Our dataset provides 6 frames of ground truth mask annotations per second, totaling 9,486 frame annotations, making it the largest VCOD dataset to date. Table 1 and Figure 4 show the characteristics of our dataset.
@@ 3.2 Dataset Statistics and Characteristics
As shown in the bottom part of Figure 1, the camouflage objects are categorized into four types, representing different subjects: animal, encompassing both wild and domesticated species that blend into their environments; human, referring to individuals who employ various techniques to remain concealed in different settings; medical target, such as objects requiring segmentation in clinical videos; and vehicle, involving transportation means like cars or tanks designed to evade detection in their respective environments.
@@ 4.2 Experiments
TSP-SAM-Point(CVPR’24) Video 0.673 0.400 0.012 0.421 0.345 0.695 0.490 0.030 0.501 0.412
TSP-SAM-Bbox(CVPR’24) Video 0.689 0.444 0.008 0.458 0.388 0.695 0.490 0.030 0.501 0.412
Ours Video 0.709 0.451 0.008 0.473 0.392 0.689 0.453 0.030 0.481 0.394
@@ 4.2.1 Implementation Details
All input images were first randomly resized to a scale ranging from 0.5 to 2 times their original size, followed by random cropping to produce 512x512 patches. The patches were then augmented using random flips and photometric distortion. The batch size was set to 12. The entire model was optimized using the AdamW [31] optimizer. In terms of the training strategy, and to ensure fairness, we followed a two-stage model training approach, similar to Cheng et al. [2]. In the first stage, the backbone was trained on the COD10k dataset, followed by fine-tuning on either the MoCA-mask or MSVCOD training set. Evaluation was then conducted on three video camouflage object detection datasets: CAD, MoCA-mask, and MSVCOD. During the pre-training phase, the learning rate was set to 1e-4 and decayed to 0 after 80,000 iterations. For fine-tuning, the learning rate was set to 1e-5, with a maximum of 10,000 training iterations.

### v2

@@ Abstract
Video Camouflaged Object Detection (VCOD) is a challenging task which aims to identify objects that seamlessly concealed within the background in videos. The dynamic properties of video enable detection of camouflaged objects through motion cues or varied perspectives. Previous VCOD datasets primarily contain animal objects, limiting the scope of research to wildlife scenarios. However, the applications of VCOD extend beyond wildlife and have significant implications in industry, art, and medical fields. Addressing this problem, we construct a new large-scale multi-domain VCOD dataset MSVCOD. To achieve high-quality annotations, we design a semi-automatic iterative annotation pipeline that reduces costs while maintaining annotation accuracy. Our MSVCOD is the largest VCOD dataset to date, introducing multiple object categories including human, animal, medical, and vehicle objects for the first time, while also expanding background diversity across various environments. This expanded scope increases the practical applicability of the VCOD task in camouflaged object detection. Alongside this dataset, we introduce a one-steam video camouflage object detection model that performs both feature extraction and information fusion without additional motion feature fusion modules. Our framework achieves state-of-the-art results on the existing VCOD animal dataset and the proposed MSVCOD. The dataset and code will be made publicly available.
@@ 1 Introduction
Camouflaged Object Detection (COD) [9] focuses on identifying and segmenting the hidden objects that closely resemble their backgrounds. Those camouflaged objects exhibit intricate visual patterns, edge breaking, texture similarities, and color matching, making them blend seamlessly into their surroundings and challenging to detect compared to traditional object detection methods [40, 54]. COD has applications in diverse fields such as medical image segmentation [11, 10], enemy detection in the search-and-rescue, art [14, 3], industry [50], scientific research [37], and agriculture [38]. Given that motion and multi-angle visual information can help to detect camouflaged objects, Video Camouflage Object Detection (VCOD) was developed. VCOD, a sub-domain of COD, focuses on segmenting out camouflage targets in the videos by utilizing motion and multi-angle visual cues.
@@
Current video camouflage datasets primarily focus on animal scenes, overlooking the diverse camouflage modalities human society. This limitation restricts the broader applicability of VCOD, particularly in fields such as medicine, search-and-rescue, and art. The lack of diverse datasets results in researchers not having enough data to train and test models on benchmark datasets. To enhance model generalization across scenes and objects, existing VCOD models [25, 53] are often pretrained on static image datasets [9]. To address this gap, we built a novel, large-scale multi-scene VCOD benchmark dataset. It includes 162 video clips across four object categories (human, animal, medical, and vehicle) and seven scenarios (aquatic, field, medical, art, jungle, desert, snowfield). Our dataset provides 6 frames of ground truth mask annotations per second, totaling 9,486 frame annotations, making it the largest VCOD dataset to date. Table 1 and Figure 4 show the characteristics of our dataset.
@@ 3.2 Dataset Statistics and Characteristics
As shown in the bottom part of Figure 1, the camouflage objects are categorized into four types, representing different subjects: animal, encompassing both wild and domesticated species that blend into their environments; human, referring to individuals who employ various techniques to remain concealed in different settings; medical target, such as objects requiring segmentation in clinical videos; and vehicle, involving transportation means like cars designed to evade detection in their respective environments.
@@ 4.2 Experiments
Ours(OSCNet) Video 0.706 0.455 0.011 0.495 0.404 0.682 0.476 0.035 0.513 0.409
@@ 4.2.1 Implementation Details
All input images were random cropping to produce 512x512 patches. The patches were then augmented using random flips and photometric distortion. The batch size was set to 8. The entire model was optimized using the AdamW [31] optimizer. In terms of the training strategy, and to ensure fairness, we followed a two-stage model training approach, similar to Cheng et al. [2]. In the first stage, the backbone was trained on the COD10k dataset, followed by fine-tuning on either the MoCA-mask or MSVCOD training set. Evaluation was then conducted on three video camouflage object detection datasets: CAD, MoCA-mask, and MSVCOD. During the pre-training phase, the learning rate was set to 1e-4 and decayed to 0 after 80,000 iterations. For fine-tuning, the learning rate was set to 1e-5, with a maximum of 8,000 training iterations.